OpenStudy (anonymous):

help me please :) 1. Which expression is equivalent to radical 5+ radical 3 / radical 5 - radical 3? 2. The fraction 1-1/a divided by 1-1/a^2 is equiv. to... 3. Solve for x: 3/x-2+8/x^2-4=1

OpenStudy (tkhunny):

Please use parentheses to write what you intend. In your #1, you have written \(\sqrt{5} + \dfrac{\sqrt{3}}{\sqrt{5}} - \sqrt{3}\). I'm guessing this is not what you intended. In what I suspect you intended, multiply numerator and denominator separetely by \(\sqrt{5}+\sqrt{3}\) and see what magic happens, especially in the denominator.

OpenStudy (anonymous):